A cave rescue team lifts an injured spelunker directly upward and out of a sinkhole by means of a motor-driven cable. The lift is performed in three stages, each requiring a vertical distance of 10.0 m: (a) the initially stationary spelunker is accelerated to a speed of 4.90 m/s; (b) he is then lifted at the constant speed of 4.90 m/s; (c) finally he is decelerated to zero speed. How much work is done on the 68.0 kg rescue by the force lifting him during each stage?
1
Expert's answer
2019-06-18T08:09:16-0400

Let's agree

"m=68kg" mass of spelunker

"s=10m" - distance for each stage

"a-" acceleration

"t-" time of movement

"g=9.8 m\/s^2"

For the srage (a):

"v_0=0"

"v=4.9m\/s" - speed in the end of stage (a)


The equation of movement for stage (a):


"v=v_0+(a-g)t""s=v_0t+(a-g)t^2\/2""\implies""a=g+v^2\/2s"

Using formula for the work made by force and Newton's second law :


"A=Fs""F=ma""A_a=m(g+v^2\/2s)s"

For the stage (b) with the constant velocity acconding the Newton's second and third laws


"F=ma=mg"

Using formula for the work made by force


"A_b=Fs=mgs"

The equation of movement for stage (c):


"v=v_0+(a-g)t""s=v_0t+(a-g)t^2\/2"

in this case "v=0" - movement until stop, "v_0=4.9m\/s" - start moving speed


"\implies""a=g-v_0^2\/2s"

Using formula for the work made by force and Newton's second law :



"A=Fs""F=ma""A_c=m(g-v_0^2\/2s)s"

The total work:


"A=A_a+A_b+A_c"

Using numbers:


"A=19992 J"

Answer: "A=19992 J"